以下にご案内の通り研究集会を行ないますのでお知らせいたします.

Asymptotic Behavior of Solutions to Partial Differential Equations

代表者 Nakao Hayashi (Osaka Univ.) Keiichi Kato (Science Univ. of Tokyo)

平成１４年２月４日ー５日

東京理科大学理窓会館２階会議室

平成１４年２月４日

10:00-10:50 E.I.Kaikina (Instituto Teconológico de Morelia)

Asymptotic behavior of solutions for some dissipative evolution equations

11:00-11:50 T.Ozawa (Hokkaido University)

Interpolation Inequalities for Besov Spaces

(lunch)

13:30-14:20 H.Takuwa (Kyoto University)

Analytic smoothing effects for a class of

dispersive equations

14:30-15:20 Y.Yamazaki (Hokkaido University)

Modified Scattering states for subcritical derivative

Nonlinear Schrödinger equations

15:30-16:20 Y.Nakamura (Kumamoto University)

The Invicsid Limit for Complex Ginzburg-Landau Equation

16:30-17:20 H.Uchida (Science University of Tokyo)

Analyticity of solutions to some nonlinear Schrödinger equations 平成１４年２月５日

10:00-10:50 P.I.Naumkin (Instituto de Matematicas, UNAM)

Survay of recent works on asymptotics of solutions to nonlinear dispaersive equations

11:00-11:50 M.Sugimito (Osaka University)

A smoothing property of Schrödinger equations and the global existence of small solutions to derivative nonlinear equations with a structure

(lunch)

13:30-14:20 H. Sunagawa (Osaka University)

Global existence and asymptotic behavior of solutions to systems of nonlinear Klein-Gordon equations with different mass terms

14:30-15:20 N. Kita (Kyushu University)

On a solution to nonlinear Schrödinger equation with δ-functional initial data

15:30-16:20 H. Takaoka (Hokkaido Uniiversity)

On global solution for the nonlinear Schrödinger equations

16:30-17:20 S.Katayama (Wakayama University)

Global existence for systems of nonlinear wave equations with different propagation speeds

なおこの研究会は文部省科学研究費研究代表者: 小澤 徹, 基盤研究 A(1) 課題番号 13304011 「波動場の幾何と解析」及び研究代表者: 林 仲夫, 基盤研究 B(2) 課題番号 12440050 「非線形分散型方程式の解の性質についての研究」の援助によりおこなわれます.